OpenStudy (anonymous):

cos^2 x + sin^2 x/cot^2 x - csc^2 x a. csc x b. -1 c. 1 d. sec x I've been having trouble with this problem for 20 minutes now and every time I get an answer it's not one of the available options to pick from. Can someone please explain this?

OpenStudy (anonymous):

the numerator is one

OpenStudy (anonymous):

\[\cos^{2}x + \sin^{2}x = 1\]

OpenStudy (anonymous):

\[\cot^{x}x - \csc^{2}x = -1\]

OpenStudy (anonymous):

and if you divide everything in the equation \[\cos^2(x)+\sin^2(x)=1\] by \[\sin^2(x)\] you will get an expression involving \(\cot^2(x)\) and \(\csc^2(x)\)

OpenStudy (anonymous):

in fact you will get \[\cot^2(x)+1=\csc^2(x)\]

OpenStudy (anonymous):

so the answer is -1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay that makes A LOT more sense now, thank you!